Parth (parthkohli):

OwlKid's intuition. Why is \(a \Large {b \over c} = {\Large ac + b \over c}?\) First intuition: \(a {\Large{b \over c}} = a + \Large{b \over c}\) Second intuition: \(a { \Large {b \over c}} = \Large{ac \over c} + \Large{b \over c}\). So, we may claim that we are multiplying \(c\) to \(a\) and getting a common denominator. Third intuition: \(\Large {ac \over c} + \Large {b \over c} = \Large{ac + b \over c}\) But if you remember, that is what we are doing! We are adding \(ac + b\) and then putting that little common denominator under! :)

OpenStudy (anonymous):

good for students who are dealing with fractions..it is reallly helpful ..

mathslover (mathslover):

students generally make mistakes .. like these

Parth (parthkohli):

Thank you, both!

OpenStudy (anonymous):

Generally we write it as: A fraction we can represent it as: \[Quotient \frac{Remainder}{Divisor}\] Suppose: 8/3 \[2\frac{2}{3}\]

Parth (parthkohli):

What's your question right there?

OpenStudy (anonymous):

I am only showing that how a fraction is converted into the form: \(a \frac{b}{c}\)

Parth (parthkohli):

Yep, but this tutorial is about something else :P

Parth (parthkohli):

I believe that it's a prerequisite for people to know that actually.

OpenStudy (anonymous):

On Dividing 8 by 3, You get Divisor = 3, Quotient = 2 and Remainder = 2.. Just put these values..

OpenStudy (anonymous):

It is obvious that: \[Dividend = Divisor \times Quotient + Remainder\]

OpenStudy (anonymous):

That is why \(a \frac{b}{c} = \frac{ac + b}{c}\)

Parth (parthkohli):

That's a good explanation too.

OpenStudy (lgbasallote):

there's one problem though... you're explaining the transformation of mixed number to improper fraction.....using algebra?

Parth (parthkohli):

Yes.

Parth (parthkohli):

People who are 13+ must know algeba